OpenStudy (pythagoras123):

Points A and B are the respective x and y-intercepts of the line y+x-c=0. Given that the length of AB is 4 units, find the positive value of c.

OpenStudy (pythagoras123):

help @shamil98 @hartnn

OpenStudy (shubhamsrg):

What is an x intercept pythagoras ?

OpenStudy (pythagoras123):

the place where the line cuts the x-axis.

OpenStudy (shubhamsrg):

Right, let us say (a,0) is the x intercept and (0,b) the y intercept then satisfy both the coordinates on the line x+y-c=0 and also use distance formula between the coordinates.

OpenStudy (pythagoras123):

i don't get it... can you show how do to it? :)

hartnn (hartnn):

to get intercepts algebraically, for x intercept, put y = 0 for y intercept, put x=0

hartnn (hartnn):

once u get 2 intercepts, use the distance formula

OpenStudy (pythagoras123):

I got \[y = -x + c \] and \[x ^{2} + y ^{2}\] = 4 but im not sure how to proceed from here cause there's 3 unknowns and only 2 equations I take that since A and B are x and y intercepts, AB is a downward sloping line - hence I used pythagoras's theorem to find AB since AB is the hypotenuse.

OpenStudy (pythagoras123):

distance formula is pythagoras theorem?

hartnn (hartnn):

x intercept ----> y= 0 0+x-c = 0 x=c so the x intercept is (c,0) got this ?

OpenStudy (pythagoras123):

yep

hartnn (hartnn):

similarly find y intercept ?

OpenStudy (pythagoras123):

but wouldn't y intercept be c also?

OpenStudy (pythagoras123):

cause in y=mx +c formula, c is the y intercept

OpenStudy (pythagoras123):

then x = y?

hartnn (hartnn):

yes so the y intercept = 0,c now u have 2 points c,0 and 0,c apply distance formula

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

hartnn (hartnn):

\(\huge 4=\sqrt{(c-0)^2+(0-c)^2}\) find c from here

OpenStudy (pythagoras123):

so c is square root 2?

hartnn (hartnn):

16 = 2c^2 c^2 = 8 how square root 2 ? wouldn't it be 2 root 2 ?

OpenStudy (pythagoras123):

oh ya forgot to square the 4

OpenStudy (pythagoras123):

thanks so much :D

hartnn (hartnn):

you're welcome ^_^